A farmer wants to fence in a rectangular pasture that is 400 yards by 224 yards. the cost of fencing is $15.75 per 8 ft. What is the total cost of the fencing?
(HINT: 3 feet = 1 yard)

Answer by math-vortex(648)

A farmer wants to fence in a rectangular pasture that is 400 yards by 224 yards. the cost of fencing is $15.75 per 8 ft. What is the total cost of the fencing?
Hi, there--
.
The farmer wants to put the fence around the edge of use rectangular pasture. This is the perimeter. Notice that this rectangular pasture has two sides that measure 400 yards and two sides that measure 224 yards. The perimeter is 400+400+224+224= 1,248 yards.
.
Since the cost of the fencing is given in dollars per feet, we need to convert 1,248 yards into feet. Every yard is equivalent to 3 feet. We have 1,248 yards, so we multiply: 1,248 x 3 = 3,744 feet.
.
Fencing costs $8 for each foot. We need to know how many 8-foots groups we have in 3,744 feet, so we divide: 3,744/8 = 468.
.
There are 468 eight-foot sections of fence. Each of these sections costs $15.75, so we multiply: 468 * 15.75 = 7,371.
.
The fencing costs $7,371.
.
Hope this helps!. Feel free to email if you have questions about this.
.
Ms.Figgy
math.in.the.vortex@gmail.com